When x number of units are sold, the price of each unit (in dollars) is given by p= -x/2+75. Find the unit price when the following quantities are sold: 2, 7, 9, 11.
Answer by checkley71(8403)

P=-X/2+75
P=-2/2+75=-1+75=$74 ANSWER.
P=-7/2+75=-3.5+75=$71.50 ANSWER.
P=-9/2+75=-4.5+75=$70.50 ANSWER
P=-11/2=75=-5.5+75=$69.50 ANSWER.